Solve the following equations:
3s = -9
Solve the following equations:
3s = -9

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

Now,
We have to divide both sides of the equation by 3,
Then, we get,
= 3s/3 = -9/3
= s = -3

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We can solve this question as follows:
3s = -9
S = -9/3
S = -3
 
Hope it helps!
Thank you ✌️